OpenStudy (anonymous):

x^2+7x=-12 what are the solutions ,

OpenStudy (anonymous):

x^2+7x+12=0 (x+3)(x+4)=0 can you solve this one?

OpenStudy (anonymous):

-3 , -4 ? is this the answer to my question ?

OpenStudy (anonymous):

that's right!

OpenStudy (anonymous):

thats not the correct answer to my equation tho

OpenStudy (anonymous):

let's check: (-3)^2 +7*(-3)+12=9-21+12=0 x=-3 is correct (satisfies your original equation).

OpenStudy (anonymous):

(-4)^2+7*(-4)+12= 16-28+12=0 this means that x=-4 is correct (satisfies your original equation).

OpenStudy (anonymous):

so, I do not know what could be wrong, unless you made a mistake in posting your original problem

OpenStudy (anonymous):

monique you dont kno wat your doin lol